OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 36x^2 − 18x − 9 4a^2 − 28a − 49 25b^2 − 40b + 16 4x^2 − 6x + 9

OpenStudy (amistre64):

id bet 3 or 4

myininaya (myininaya):

\[(6x-3)^2=36x^2-2(6)(3)x+9 \] \[(2a-7)^2=4a^2-2(2)(7)+49\] \[(5b-4)^2=25b^2-2(5)(4)+16\] \[(2x-3)^2=4x^2-2(2)(3)+9\]

OpenStudy (amistre64):

ohohohoh..... discriminate of zero lol

myininaya (myininaya):

-2(2)(3) is not -6 so whats left? i give you a hint look at choice 3

OpenStudy (amistre64):

25b^2 − 40b + 16 40^2 -4(25)(16) 1600 -100(16) 1600 -1600 = 0

OpenStudy (amistre64):

.... also? ;)

OpenStudy (amistre64):

i figured it needed a + on the rearend

OpenStudy (amistre64):

and 4 just looked rather weak and pathetic

myininaya (myininaya):

yes it does

OpenStudy (anonymous):

What are the factors of 3x^2 + 10x + 8?